OpenStudy (anonymous):

find prime (x) for f(x)=3^(2x+6)

OpenStudy (anonymous):

Sorry does this ask when will f(x) be prime?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

x is all the real numbers?

OpenStudy (anonymous):

oh yes sorry i didnt look close at question. yes

OpenStudy (anonymous):

but than there are infinite many solutions as your f(x) is continuous and takes all values from 0 till infinity

OpenStudy (anonymous):

Isnt the question to find x prime as the derivative?

OpenStudy (anonymous):

you re looking for the derivative? answer is \[f'(x)=2\ln(3)\times 3^{2x+6}\]by the chain rule

OpenStudy (anonymous):

\[\frac{d}{dx}a^x=2^x\ln(a)\]